How would you write 10 to the -6 in decimal form? Also, another problem, 6x10 to the -3? If you can send me the answer, and how you got that, as SOON as possible, like later on today, or early-bout 7 or 8 in the morning, I would TRULY AND GLADLY appreciate it!


2/13/12

Answer by scott8148(6628)

10^2 = 100
10^1 = 10
10^0 = 1
10^-1 =.1
10^-2 = .01

follow the pattern to 10^-6

6x10^-3 is just what it says , 6 multiplied by 10^-3